OpenStudy (anonymous):

help? a man drives his car at a uniform rate from his home toward albany ny. the man;s son left the same house 45 minutes later than his father and traveled along the same road at a rate 12miles per hour faster than his father. the son overtook his father 130 miles from home. what was the rate at which the father drove?

OpenStudy (anonymous):

total time that took son to reach 130miles = 130/12=10.833 total time that took father to reach 130miles = 10.833 + 0.75 = 11.5833 hours (here 0.75hours=45mins) rate at which father drove = 130/11.5833 = 11.223 miles per hours

OpenStudy (anonymous):

ok, but i have an answer sheet that has a write answer and a wrong answer and u have to choose the right one, the two choices are 40 and 2.4

OpenStudy (anonymous):

well to my knowledge, I think I got the right answer. If you really want to stick to one of your answers, then think like this. father's rate should be lower than son's because it is obvious here. So 40 cannot be the answer.

OpenStudy (anonymous):

ok thank you for your help